OpenStudy (anonymous):

Use the addition formula to derive the identity: cos(x-pi/2) = sinx

OpenStudy (anonymous):

use \[\cos(\alpha - \beta)=\cos(\alpha)\cos(\beta)+\sin(\alpha)\sin(\beta)\] with \[\alpha = x, \beta = \frac{\pi}{2}\]

OpenStudy (anonymous):

get \[\cos(x - \frac{\pi}{2})=\cos(x)\cos(\frac{\pi}{2})+\sin(x)\sin(\frac{\pi}{2})\] and then since \[\cos(\frac{\pi}{2})=0\] and \[\sin(\frac{\pi}{2})=1\] you get your result

OpenStudy (anonymous):

you're my hero